# Measurement of the branching fraction of $D^{+}_{s}\to\ell^{+}\nu_{\ell}$ via $e^{+}e^{-}\to D^{*+}_{s}D^{*-}_{s}$

### v2

@@ Abstract
Based on $10.64~{}\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ collision data taken at center-of-mass energies between 4.237 and 4.699 GeV with the BESIII detector, we study the leptonic $D^{+}_{s}$ decays using the $e^{+}e^{-}\to D^{*+}_{s}D^{*-}_{s}$ process. The branching fractions of $D_{s}^{+}\to\ell^{+}\nu_{\ell}\,(\ell=\mu,\tau)$ are measured to be $\mathcal{B}(D_{s}^{+}\to\mu^{+}\nu_{\mu})=(0.547\pm 0.026_{\rm stat}\pm 0.016_{\rm syst})\%$ and $\mathcal{B}(D_{s}^{+}\to\tau^{+}\nu_{\tau})=(5.60\pm 0.16_{\rm stat}\pm 0.20_{\rm syst})\%$, respectively. The product of the decay constant and Cabibbo-Kobayashi-Maskawa matrix element $|V_{cs}|$ is determined to be $f_{D_{s}^{+}}|V_{cs}|=(246.5\pm 5.9_{\rm stat}\pm 3.6_{\rm syst}\pm 0.5_{\rm input})_{\mu\nu}~{}\mathrm{MeV}$ and $f_{D_{s}^{+}}|V_{cs}|=(252.7\pm 3.6_{\rm stat}\pm 4.5_{\rm syst}\pm 0.6_{\rm input}))_{\tau\nu}~{}\mathrm{MeV}$, respectively. Taking the value of $|V_{cs}|$ from a global fit in the Standard Model, we obtain ${f_{D^{+}_{s}}}=(252.8\pm 6.0_{\rm stat}\pm 3.7_{\rm syst}\pm 0.6_{\rm input})_{\mu\nu}$ MeV and ${f_{D^{+}_{s}}}=(259.2\pm 3.6_{\rm stat}\pm 4.5_{\rm syst}\pm 0.6_{\rm input})_{\tau\nu}$ MeV, respectively. Conversely, taking the value for $f_{D_{s}^{+}}$ from the latest lattice quantum chromodynamics calculation, we obtain $|V_{cs}|=(0.986\pm 0.023_{\rm stat}\pm 0.014_{\rm syst}\pm 0.003_{\rm input})_{\mu\nu}$ and $|V_{cs}|=(1.011\pm 0.014_{\rm stat}\pm 0.018_{\rm syst}\pm 0.003_{\rm input})_{\tau\nu}$, respectively.
@@ V.1 Selection of double-tag events
The candidates for $D^{*+}_{s}$ containing a leptonic $D^{+}_{s}$ decay are reconstructed from the unused showers and tracks remaining after the single-tag selection. We require only one additional track not used in the tag reconstruction: $N_{\rm extra}^{\rm charge}$= 0. The transition $\gamma(\pi^{0})$ is selected with the same selection criteria as the tag side.
Throughout this paper, the $\mu_{a}$ denote the $\mu^{+}$ candidates identified with MUC information, the $\mu_{b}$ denote the $\mu^{+}$ candidates identified without MUC information; and the $\tau^{+}_{e}$, $\tau^{+}_{\mu}$, $\tau^{+}_{\pi}$, and $\tau^{+}_{\rho}$ denote the $\tau^{+}$ candidates reconstructed via $\tau^{+}\to e^{+}\nu_{e}\bar{\nu}_{\tau}$, $\tau^{+}\to\mu^{+}\nu_{\mu}\bar{\nu}_{\tau}$, $\tau^{+}\to\pi^{+}\bar{\nu}_{\tau}$, and $\tau^{+}\to\rho^{+}\bar{\nu}_{\tau}$, respectively. In the selection of the candidates for $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$, the $\pi^{+}$ and $\rho^{+}$ candidates are selected with the same selection criteria as those used on the tag side. To select the candidates for $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, the positron PID uses the measured information in the MDC, TOF, and EMC. The combined likelihoods ($\mathcal{L}^{\prime}$) under the positron, pion, and kaon hypotheses are obtained. Positron candidates are required to satisfy $\mathcal{L}^{\prime}(e)>0.001$ and $\mathcal{L}^{\prime}(e)/(\mathcal{L}^{\prime}(e)+\mathcal{L}^{\prime}(\pi)+\mathcal{L}^{\prime}(K))>0.8$. To select the candidates for $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, the muon candidate is required to have a deposited energy in the EMC within (0.1, 0.3)0.10.3(0.1,\,0.3)( 0.1 , 0.3 ) GeV. It must also satisfy a requirement on the hit depth, $d_{\mu^{+}}$, in the muon counter which depends on both $|\cos\theta_{\mu^{+}}|$ and $p_{\mu^{+}}$. These requirements on $d_{\mu^{+}}$ are shown in Table 3. The selected $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ candidates offer an opportunity to determine the BF of $D^{+}_{s}\to\mu^{+}_{b}\nu_{\mu}$ as a cross check.
@@
Figure 2 shows the $M^{2}_{\rm miss}$ distribution of the candidate events for $D^{+}_{s}\to\mu_{a}^{+}\nu_{\mu}$. The left two plots of Fig. 3 show the $E_{\rm sum}^{\rm extra\,\gamma}$ distributions of the candidate events for $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, and the right two plots of Fig. 3 show the $M^{2}_{\rm miss}$ distributions of the candidate events for $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$. Different signal variables are chosen for the different $\tau^{+}$ modes mainly to achieve better separation between signal and background. These distributions are obtained from the combined data from all energy points.
@@ V.2 Detection efficiencies
Figure 3: Simultaneous fit to the $E_{\rm sum}^{\rm extra~{}\gamma}$ distributions (left) of $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$ as well as the $M_{\rm miss}^{2}$ distributions (right) of $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$. The points with error bars represent data combined from all energy points. The blue solid curve denotes the total fit. The red solid curves and red dashed curves show the fitted combinatorial background shapes. The pink dashed, cyan dashed and blue dashed are the backgrounds from $D_{s}^{+}\to\mu^{+}\nu_{\mu^{+}}$, $D_{s}^{+}\to K^{0}\pi^{+}$, $D_{s}^{+}\to\eta\pi^{+}\pi^{0}$ and $D_{s}^{+}\to K^{0}\pi^{+}\pi^{0}$, respectively.
@@ V.3 Branching fractions without lepton universality constraint
To extract the BF of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$, we perform a simultaneous fit to the $E_{\rm sum}^{\rm extra~{}\gamma}$ distributions of candidates for $D^{+}_{s}\to\tau^{+}_{\ell}\tau_{\tau}$ ($\ell=e$ or $\mu$) and the $M_{\rm miss}^{2}$ distributions of candidates for $D^{+}_{s}\to\tau^{+}_{h}\tau_{\tau}$,( $h=\pi^{+}$ or $\rho^{+}$), as shown in Fig. 3. There is large component of $D^{+}_{s}\to\mu_{b}^{+}\nu_{\mu}$ in the accepted candidates $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ due to the poorer separation muons and pions when using only the $dE/dx$, TOF, and EMC information. The signal shapes of $D^{+}_{s}\to\mu_{a}^{+}\nu_{\mu}$, $D^{+}_{s}\to\mu_{b}^{+}\nu_{\mu}$, $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$ are modeled by the individual simulated shapes for each decay mode. The signal shapes of the $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$ are described by a sum of two bifurcated-Gaussian functions, whose parameters are determined from the fits to the signal MC events and are fixed in the simultaneous fit. The background components for each signal decay are modeled with shapes derived from the relevant MC simulation The four $\tau^{+}$ decay modes are constrained to have a common BF for $D^{+}_{s}\to\tau^{+}\nu_{\tau}$, taking into account different efficiencies and daughter particle decay BFs. From this simultaneous fit, we obtain
@@ VIII Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2020YFA0406300, 2020YFA0406400, 2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 11961141012, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.

### v3

@@ Abstract
Based on $10.64~{}\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ collision data taken at center-of-mass energies between 4.237 and 4.699 GeV with the BESIII detector, we study the leptonic $D^{+}_{s}$ decays using the $e^{+}e^{-}\to D^{*+}_{s}D^{*-}_{s}$ process. The branching fractions of $D_{s}^{+}\to\ell^{+}\nu_{\ell}\,(\ell=\mu,\tau)$ are measured to be $\mathcal{B}(D_{s}^{+}\to\mu^{+}\nu_{\mu})=(0.547\pm 0.026_{\rm stat}\pm 0.016_{\rm syst})\%$ and $\mathcal{B}(D_{s}^{+}\to\tau^{+}\nu_{\tau})=(5.60\pm 0.16_{\rm stat}\pm 0.20_{\rm syst})\%$, respectively. The product of the decay constant and Cabibbo-Kobayashi-Maskawa matrix element $|V_{cs}|$ is determined to be $f_{D_{s}^{+}}|V_{cs}|=(246.5\pm 5.9_{\rm stat}\pm 3.6_{\rm syst}\pm 0.5_{\rm input})_{\mu\nu}~{}\mathrm{MeV}$ and $f_{D_{s}^{+}}|V_{cs}|=(252.7\pm 3.6_{\rm stat}\pm 4.5_{\rm syst}\pm 0.6_{\rm input}))_{\tau\nu}~{}\mathrm{MeV}$, respectively. Taking the value of $|V_{cs}|$ from a global fit in the Standard Model, we obtain ${f_{D^{+}_{s}}}=(253.2\pm 6.0_{\rm stat}\pm 3.7_{\rm syst}\pm 0.6_{\rm input})_{\mu\nu}$ MeV and ${f_{D^{+}_{s}}}=(259.6\pm 3.7_{\rm stat}\pm 4.6_{\rm syst}\pm 0.6_{\rm input})_{\tau\nu}$ MeV, respectively. Conversely, taking the value for $f_{D_{s}^{+}}$ from the latest lattice quantum chromodynamics calculation, we obtain $|V_{cs}|=(0.986\pm 0.023_{\rm stat}\pm 0.014_{\rm syst}\pm 0.003_{\rm input})_{\mu\nu}$ and $|V_{cs}|=(1.011\pm 0.014_{\rm stat}\pm 0.018_{\rm syst}\pm 0.003_{\rm input})_{\tau\nu}$, respectively.
@@ V.1 Selection of double-tag events
The candidates for $D^{*+}_{s}$ containing a leptonic $D^{+}_{s}$ decay are reconstructed from the unused showers and tracks remaining after the single-tag selection. We require that only one additional track remains after the tag reconstruction. To further suppress hadronic background, we require that there is no extra charged track in each candidate event: $N_{\rm extra}^{\rm charge}$= 0. The transition $\gamma(\pi^{0})$ is selected with the same selection criteria as the tag side.
Throughout this paper, the $\mu^{+}_{a}$ denote the $\mu^{+}$ candidates identified with MUC information, the $\mu_{b}$ denote the $\mu^{+}$ candidates identified without MUC information; and the $\tau^{+}_{e}$, $\tau^{+}_{\mu}$, $\tau^{+}_{\pi}$, and $\tau^{+}_{\rho}$ denote the $\tau^{+}$ candidates reconstructed via $\tau^{+}\to e^{+}\nu_{e}\bar{\nu}_{\tau}$, $\tau^{+}\to\mu^{+}\nu_{\mu}\bar{\nu}_{\tau}$, $\tau^{+}\to\pi^{+}\bar{\nu}_{\tau}$, and $\tau^{+}\to\rho^{+}\bar{\nu}_{\tau}$, respectively. In the selection of the candidates for $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$, the $\pi^{+}$ and $\rho^{+}$ candidates are selected with the same selection criteria as those used on the tag side. To select the candidates for $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, the positron PID uses the measured information in the MDC, TOF, and EMC. The combined likelihoods ($\mathcal{L}^{\prime}$) under the positron, pion, and kaon hypotheses are obtained. Positron candidates are required to satisfy $\mathcal{L}^{\prime}(e)>0.001$ and $\mathcal{L}^{\prime}(e)/(\mathcal{L}^{\prime}(e)+\mathcal{L}^{\prime}(\pi)+\mathcal{L}^{\prime}(K))>0.8$. To select the candidates for $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, the muon candidate is required to have a deposited energy in the EMC within (0.1, 0.3)0.10.3(0.1,\,0.3)( 0.1 , 0.3 ) GeV. It must also satisfy a requirement on the hit depth, $d_{\mu^{+}}$, in the muon counter which depends on both $|\cos\theta_{\mu^{+}}|$ and $p_{\mu^{+}}$. These requirements on $d_{\mu^{+}}$ are shown in Table 3. The selected $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ candidates offer an opportunity to determine the BF of $D^{+}_{s}\to\mu^{+}_{b}\nu_{\mu}$ as a cross check.
@@
Figure 2 shows the $M^{2}_{\rm miss}$ distribution of the candidate events for $D^{+}_{s}\to\mu_{a}^{+}\nu_{\mu}$. The left two plots of Fig. 3 show the $E_{\rm sum}^{\rm extra\,\gamma}$ distributions of the candidate events for $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, and the right two plots of Fig. 3 show the $M^{2}_{\rm miss}$ distributions of the candidate events for $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$. Different signal variables are chosen for the different $\tau^{+}$ modes mainly to achieve better separation between signal and background. There are no prominent peaking backgrounds for $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, while the main peaking backgrounds are from $D^{+}_{s}\to\mu^{+}\nu_{\mu}$ and $D^{+}_{s}\to K^{0}\pi^{+}$ for $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$, and the main peaking backgrounds are from $D^{+}_{s}\to\eta\pi^{+}\pi^{0}$ and $D^{+}_{s}\to K^{0}\pi^{+}\pi^{0}$ for $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$. These distributions are obtained from the combined data from all energy points.
@@ V.2 Detection efficiencies
Figure 3: Simultaneous fit to the $E_{\rm sum}^{\rm extra~{}\gamma}$ distributions (left) of $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$ as well as the $M_{\rm miss}^{2}$ distributions (right) of $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$. The points with error bars represent data combined from all energy points. The blue solid curve denotes the total fit. The red solid curves and red dashed curves show the fitted combinatorial background shapes. The pink dashed, cyan dashed blue dashed, and green dashed are the backgrounds from $D_{s}^{+}\to\mu^{+}\nu_{\mu^{+}}$, $D_{s}^{+}\to K^{0}\pi^{+}$, $D_{s}^{+}\to\eta\pi^{+}\pi^{0}$ and $D_{s}^{+}\to K^{0}\pi^{+}\pi^{0}$, respectively.
@@ V.3 Branching fractions without lepton universality constraint
To extract the BF of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$, we perform a simultaneous fit to the $E_{\rm sum}^{\rm extra~{}\gamma}$ distributions of candidates for $D^{+}_{s}\to\tau^{+}_{\ell}\tau_{\tau}$ ($\ell=e$ or $\mu$) and the $M_{\rm miss}^{2}$ distributions of candidates for $D^{+}_{s}\to\tau^{+}_{h}\tau_{\tau}$,( $h=\pi^{+}$ or $\rho^{+}$), as shown in Fig. 3. There is large component of $D^{+}_{s}\to\mu_{b}^{+}\nu_{\mu}$ in the accepted candidates $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ due to the poorer separation muons and pions when using only the $dE/dx$, TOF, and EMC information. The signal shapes of $D^{+}_{s}\to\mu_{a}^{+}\nu_{\mu}$, $D^{+}_{s}\to\mu_{b}^{+}\nu_{\mu}$, $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$ are modeled by the individual simulated shapes for each decay mode. The signal shapes of the $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$ are described by a sum of two bifurcated-Gaussian functions, whose parameters are determined from the fits to the signal MC events and are fixed in the simultaneous fit. The background components for each signal decay are modeled with shapes derived from the relevant MC simulation. The four $\tau^{+}$ decay modes are constrained to have a common BF for $D^{+}_{s}\to\tau^{+}\nu_{\tau}$, taking into account different efficiencies and daughter particle decay BFs. From this simultaneous fit, we obtain
@@ VIII Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2023YFA1606704, 2020YFA0406300, 2020YFA0406400; National Natural Science Foundation of China (NSFC) under Contracts Nos. 12375092, 11635010, 11735014, 11935015, 11935016, 11935018, 11961141012, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.